OpenStudy (anonymous):

Verify the identity. cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x

OpenStudy (abb0t):

What are you being axed 2 do?

OpenStudy (anonymous):

Verify the identity

OpenStudy (abb0t):

You can factor out a 2 from the right function first of all. also, note that \[\sin^2(x) = \cos(x)-\]

OpenStudy (abb0t):

\[\cos^2(x)-1\]

OpenStudy (anonymous):

=2(1-1 cos^2-1 x - 1 cos^2-1x)

OpenStudy (anonymous):

what should I do next?

OpenStudy (abb0t):

Wait, what is ur function, I'm trying to see what you wrote, correct me if I'm wrong, but is this your function: \[\cos(4x)+\cos(2x) = 2-2\sin^2(2x)-2\sin^2(x)\]

OpenStudy (anonymous):

thats correct